### v2

# Open-Source Molecular Processing Pipeline for Generating Molecules

Abstract
Generative models for molecules have shown considerable promise for use in computational chemistry, but remain difficult to use for non-experts. For this reason, we introduce open-source infrastructure for easily building generative molecular models into the widely used DeepChem (Ramsundar et al., 2019) library with the aim of creating a robust and reusable molecular generation pipeline. In particular, we add high quality PyTorch (Paszke et al., 2019) implementations of the Molecular Generative Adversarial Networks (MolGAN)  (Cao and Kipf, 2022) and Normalizing Flows (Papamakarios et al., 2021). Our implementations show strong performance comparable with past work (Kuznetsov and Polykovskiy, 2021; Cao and Kipf, 2022).
Keywords Machine Learning  $\cdot$ open-source  $\cdot$ deepchem  $\cdot$ molecule generation  $\cdot$ molecules  $\cdot$ drug discovery
1 Introduction
The discovery of new molecules and materials is crucial for addressing challenges in chemistry, such as treating diseases and tackling climate change (Liu et al., 2023; Sanchez and Aspuru-Guzik, 2018). Traditional methods rely on human expertise and are time-consuming and costly, limiting the exploration of the vast chemical space (Polishchuk et al., 2013). Generative models offer a promising solution using deep learning to design molecules based on desired properties, rapidly identifying diverse and optimized molecules for specific applications. These models vary in their approaches and have seen rapid development, with benchmarks now in place to evaluate their performance in terms of distribution learning and chemical diversity. Although these models are publicly available, practitioners require extensive Python and machine learning knowledge to reap their benefits. Thus, we introduce open-source molecular generative model infrastructure into DeepChem Ramsundar et al. (2019), a widely used open-source library for molecular machine learning.
Among these models, Molecular Generative Adversarial Networks (MolGAN) (Cao and Kipf, 2022) and Normalizing Flow models (Papamakarios et al., 2021) offer two complementary approaches to the computational generation of molecular structures. These models, though available openly, require the practitioners to have a deep understanding of machine learning and Python, which presents significant challenges and hinders their practical adoption
MolGAN adapts the generative adversarial network (Goodfellow et al., 2014) framework to molecular contexts, utilizing an adversarial training mechanism to facilitate the generation of novel molecular configurations by approximating the distribution of empirical molecular data. Normalizing Flow models use an alternative strategy through exact likelihood models, employing diffeo-morphisms to convert simple distributional forms into complex molecular distributions.
We contribute implementations of MolGAN and Normalizing flow models in PyTorch to DeepChem and build example pipelines for training generative models. This adaptation broadens the accessibility of MolGAN and Normalizing Flows to the computational chemistry community. Our pipelines require little previous knowledge in machine learning or Python to generate molecules. On the other hand, the implementations are well-documented for advanced users who want to tinker around and make modifications as per their needs.
2 Methods and Background
2.1 Deepchem and Generative Molecular Models
DeepChem (Ramsundar et al., 2019) is a highly versatile open-source Python library designed specifically for scientific machine learning and deep learning on molecular and quantum datasets (Ramsundar et al., 2021). Its comprehensive framework enables the solution of complex scientific problems in diverse areas such as drug discovery, energy calculations, and biotech (Wu et al., 2018). DeepChem’s design philosophy is centered around a systematic approach whereby scientific calculations are broken down into workflows built from underlying primitives. This approach enables the DeepChem library to apply to a wide range of applications. Notably, DeepChem has facilitated large-scale benchmarking for molecular machine learning via the MoleculeNet benchmark suite (Wu et al., 2018), protein-ligand interaction modelling (Gomes et al., 2017), generative modeling of molecules (Frey et al., 2023), and other cutting-edge scientific endeavors.
DeepChem had older TensorFlow implementations of MolGANs and Normalizing Flows. However, DeepChem has been migrating to PyTorch following the adoption of PyTorch into the Linux Foundation core for long-term stability. In this work, we migrate the older code to PyTorch and integrate the generative framework more tightly with DeepChem’s ecosystem. The implementation allows users to experiment with the vast library of existing layers that can be plugged into the models to make new and creative models.
2.2 Representation of Molecules
The strength of neural networks lies in their ability to take in a complex input representation and transform it into a latent representation needed to solve a particular task. In this way, the choice of input representation plays a key role in governing how the model learns information about the molecule. Input representations often fall into one of two categories: (1) one-dimensional (e.g., string-based representations), (2) two-dimensional (e.g., molecular graphs).
2.2.1 One-dimensional representations
The most common one-dimensional representation of molecules is SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988), which transforms a molecule into a sequence of characters based on predefined atom ordering rules. This representation enables the use of neural network architectures developed for language processing. For instance, previous work (Gómez-Bombarelli et al., 2018; Olivecrona et al., 2017) used recurrent neural networks as generative models to create SMILES strings. However, these methods often produce invalid SMILES that cannot be converted to molecular structures due to their disregard for SMILES grammar.  Kusner et al. (2017) and  Dai et al. (2018) attempted to address this by incorporating syntactic constraints into the recurrent neural network, yet they still frequently generated invalid SMILES strings. To overcome this limitation, Krenn et al. (2020) introduced SELFIES (Self-Referencing Embedded Strings), an improved string representation. With SELFIES, a recurrent neural network can generate molecules with 100% validity, though validity here pertains to valency rules and does not guarantee molecular stability.
2.2.2 Molecules as Graphs
Molecules can alternatively be represented as graphs. Graphical representations are powerful in that they directly capture the connectivity between atoms, while in one-dimensional representations, this information must inferred by the model. For MolGAN, each molecule can be represented as an undirected Graph $G$ with a set of edges $E$ and nodes $V$. Each atom $v_{i}\in V$ is associated with a $D$-dimensional one-hot vector $\mathbf{x}_{i}$. Each edge $(v_{i},v_{j})\in E$ is also associated with a bond type $y\in\{1,\dots,Y\}$. Thus, we have a representation of a graph as two objects: a $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]^{T}\in\mathbb{R}^{N\times D}$ and an adjacency tensor $\mathbf{A}\in\mathbb{R}^{N\times N\times Y}$. $\mathbf{A}_{i,j}\in\mathbb{R}^{Y}$ is a one-hot vector indicating the type of edge between $i$ and $j$.
2.3 MolGAN
Molecular Generative Adversarial Network (MolGAN) represents a novel approach in the Generative Space for molecules by employing the GAN framework, which allows for an implicit, likelihood-free generative model which overcomes various problems (like graph matching (Simonovsky and Komodakis, 2018) and node ordering heuristics) with previous models. MolGAN performs similarly to current SMILES-based approaches as well, albeit it is more susceptible to model collapse. We have incorporated the following loss used by the paper
$L(\mathbf{x}^{(i)},G_{\theta}(z^{(i)});\phi)=\underbrace{-D_{\phi}(\mathbf{x}^{(i)})+D_{\phi}(G_{\theta}(z^{(i)}))}_{\text{Original WGAN Loss}}+\\ \underbrace{\alpha(\lvert\lvert\nabla_{\hat{x}^{(i)}}D_{\phi}(\hat{x}^{(i)})\rvert\rvert-1)^{2}}_{\text{Gradient Penalty}}$ (1)
This is an improved from of the WGAN (Arjovsky et al., 2017) loss, where $G_{\theta}$ is the generator, $D_{\phi}$ is the discriminator, $\mathbf{x}^{(i)}\sim p_{data}(\mathbf{x})$ and $\mathbf{z}^{(i)}\sim p_{z}(\mathbf{z})$ and $\mathbf{\hat{x}}^{(i)}$ is a sampled linear combination as the following equation $\mathbf{\hat{x}}^{(i)}=\epsilon\mathbf{x}^{(i)}+(1-\epsilon)G_{\theta}(\mathbf{z}^{(i)})$ with $\epsilon\sim U(0,1)$. The model architecture is displayed in Figure 1.
Figure 1: Model Architecture of MolGAN
2.4 Normalizing Flows
Normalizing Flows constitute a generative model that uses invertible transformations to model a probability distribution. This methodology enables the computation of likelihoods and the generation of samples by transforming simple base distributions (e.g., gaussian) into more complex ones through a sequence of invertible (& differentiable) transformations. The precision and control afforded by Normalizing Flows make them particularly advantageous for generating molecules with highly specific attributes. Normalizing flows also take advantage of equivariance (through invariance) in the Graph’s probability distribution, thus providing a complementary strength to adversarial approaches such as MolGAN in the generative modeling landscape. The model architecture is displayed in Figure 2.
Figure 2: Model Architecture of Normalizing Flow models
2.5 Standardizing Molecule Generation Workflows
We aim to standardize (“deepchemize”  (Ramsundar et al., 2021)) the molecule generation by MolGAN and Normalizing Flow model by implementing the models using the workflow structure implemented by other DeepChem models. With this standardization, we aim to experimenting with new generative models simpler and enable better molecule generation capabilities.
BasicMolGANModel (the MolGAN implementation in DeepChem), can be used with different discriminators and generators, thus providing a flexible framework for molecule generation. This flexibility of the BasicMolGANModel can be used to incorporate more complex architectures of generators and discriminators to achieve better global density information to approximate a more accurate exchange-correlation functional. We allow for a “plug-and-play" style of training, inference, and generation with respect to the MPNN used, optimizer, number of MLP layers in the generator, and bond/node types. This can be used in a creatively to combine the strengths of multiple architectures.
NormalizingFlowModel (the Normalizing Flow implementation in DeepChem) can be used with different flow layers with customized bijective transformations. This allows end users to incorporate more complex transformations and modeling capabilities into the framework. The interoperable nature of DeepChem’s layers enables further innovation. Overall, these abstractions simplify and ease usage and experimentation to achieve state-of-the-art molecule generation.
We have also created pedagogical tutorials for all layers and networks introduced to provide a streamlined pipeline for end-users, as well as, enable fast experimentation as the mutable parameters are clearly demarcated.
3 Implementation
The Generative models were split and implemented as the following major components: Layers, Base Model, and Molecule Generation Pipeline. We briefly also describe the DeepChem Datasets used.
3.1 Datasets
Our work utilizes three publicly available datasets: QM7 (Rupp et al., 2012), BBBP (Martins et al., 2012), Lipophilicity (Hersey, 2015), PPB (Wang et al., 2005) and QM9 (Ramakrishnan et al., 2014). QM7 is a subset of GDB-13 (a database of nearly 1 billion stable and synthetically accessible organic molecules) containing up to 7 heavy atoms C, N, O, and S. The blood-brain barrier penetration (BBBP) dataset encompasses approximately 2,000 molecules with binary labels on their permeability properties. The lipophilicity dataset, curated from ChEMBL database, contains 4200 compounds. The PPB dataset is from the Maximum Unbiased Validation (MUV) group, it is a benchmark dataset selected from PubChem BioAssay by applying a refined nearest neighbor analysis. It contains around eleven thousand compounds. QM9 is a subset of the GDB-17 database (a database of nearly 166 billion stable and synthetically accessible organic molecules), comprising 134 thousand stable organic molecules with up to 9 heavy atoms. All the mentioned datasets are downloaded through MoleculeNet (Wu et al., 2018).
3.2 Layers
MolGAN: The first layer in the Discriminator of a MolGAN model is an encoder layer, which consists of multiple convolutional layers followed by an aggregation layer. We provide an architecture that can be varied and overridden easily based on the end user’s requirements.
Normalizing Flows: Layers in Normalizing Flows are used to perform forward and backward computation. DeepChem Normalizing Flows pipeline supports linear layers, which offer low computational costs compared with other layers (e.g. planar, autoregressive). Linear transformations exhibit relationships through molecular data dimensions.
$\mathbf{g(x)}=\mathbf{Wx}+\mathbf{b}$
Where $W\in\mathbb{R}^{D\times D}$ and $b\in\mathbb{R}^{D}$ are parameters. So, if W is an invertible matrix, the function is invertible.
3.3 Base Model
The Base Model has three distinct components, which can be modified individually to enable flexibility.
[…]
$z\sim\mathcal{N}(\mathbf{0,I})$
the generator $\mathcal{G}_{\theta}(z)$ outputs two continuous and dense objects,
$\mathbf{X}\in\mathbb{R}^{N\times D},\mathbf{A}\in\mathbb{R}^{N\times N\times D}$
Over which we perform categorical sampling using the Gumbel Softmax trick (Jang et al., 2017) to obtain $\mathbf{\tilde{X}}$ and $\mathbf{\tilde{A}}$ as -
$\mathbf{\tilde{X}}$ $=\mathbf{X}+\text{Gumbel}(\mu=0,\beta=1)$ (2)
$\mathbf{\tilde{A}}$ $=\mathbf{A}_{ijy}+\text{Gumbel}(\mu=0,\beta=1)$ (3)
This method can be overridden to use a straight-through estimator (Paulus et al., 2020) or can also allow for $\mathbf{\tilde{X}}=\mathbf{X}$ and $\mathbf{\tilde{A}}=\mathbf{A}$.
Discriminator - The discriminator is tasked with scoring the graph produced. This involves a series of graph-convolutional or message-passing layers followed by MLPs; the equations are as follows:
$\boldsymbol{h}_{i}^{\prime(\ell+1)}$ $=f_{s}^{(\ell)}\left(\boldsymbol{h}_{i}^{(\ell)},\boldsymbol{x}_{i}\right)+\sum_{j=1}^{N}\sum_{y=1}^{Y}\frac{\tilde{\boldsymbol{A}}_{ijy}}{\left|\mathcal{N}_{i}\right|}f_{y}^{(\ell)}\left(\boldsymbol{h}_{j}^{(\ell)},\boldsymbol{x}_{j}\right)$ (4)
$\boldsymbol{h}_{i}^{(\ell+1)}$ $=\tanh\left(\boldsymbol{h}_{i}^{\prime(\ell+1)}\right)$ (5)
We employ Relational GCN (Schlichtkrull et al., 2017) layers, which allow us to model the graph structure with multiple edge types, this shows the update of node representations per layer. Here, $h_{i}^{(l)}$ represents the signal of node $i$ at layer $l$ and $f_{s}^{(l)}$ represents a self-connection linear transformation. $\mathcal{N}_{i}$ represents the neighbours of node $i$.
$\boldsymbol{h}_{\mathcal{G}}^{\prime}$ $=\sum_{v\in\mathcal{V}}\sigma\left(i\left(\boldsymbol{h}_{v}^{(L)},\boldsymbol{x}_{v}\right)\right)\odot\tanh\left(j\left(\boldsymbol{h}_{v}^{(L)},\boldsymbol{x}_{v}\right)\right)$ (6)
$\boldsymbol{h}_{\mathcal{G}}$ $=\tanh(\boldsymbol{h}_{\mathcal{G}}^{\prime})$ (7)
After a few layers of graph convolution, we use two MLPs $i$ and $j$ similar to  (Li et al., 2017) to extract the overall graph score. $\sigma$ stands for the sigmoid function $\sigma=\frac{1}{1+e^{-x}}$.
Normalizing Flow model starts with a base distribution $p_{z}$ and performs a series of layer transformations to estimate a density function $p_{x}$, i.e.,
$x=L(z)$
where $z\sim p_{z}$, importantly, these transformations must be invertible; and so,
$\rho_{X}(x)=\rho_{Z}(z)\cdot|det(J_{L}(z))|^{-1}$
Jacobian determinant, the second term on the right side of the equation, updates the direction of the flow based on linear transformations.
3.4 Molecule Generation Pipelines
3.4.1 MolGAN
Figure 3: Molecule generation pipeline for MolGAN
Training / Generating from MolGAN on custom datasets follows a very simple pipeline, which can be achieved in a few lines of code; we describe the pipeline in Figure 3, which involves extracting SMILES (Weininger, 1988) representations and wrapping them in a DeepChem dataset that can be used to train the model.
gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
The full pipeline is in Appendix A.
3.4.2 Normalizing Flows
Figure 4: Molecule generation pipeline for Normalizing flows
Analogous to the MolGAN pipeline, the Normalizing Flows pipeline (Figure 4) expects a SELFIES (Krenn et al., 2020) string as input. There are added steps for Dequantization (i.e., adding noise in [0, 1) to every input) and test-train-valid splitter (unlike MolGAN, where you train on the full dataset).
1
2nfm = NormFlow(flows, distribution, dim)
3
4nfm.fit(max_iterations, optimizer)
5
The full pipeline is in Appendix B.
4 Experimental Results
4.1 Experimental Setup
We report results averaged across five runs of 1000 molecule generations, matching Cao and Kipf (2022).
For MolGAN evaluation, we allowed four edge types (single, double, triple, no-bond, and aromatic), five node types (C, N, O, F, and PAD), and a maximum of nine vertices per graph; this is borrowed from the original implementation along with a learning rate of $10^{-3}$. However, we use an embedding dimension of 100, no dropout (Srivastava et al., 2014), and a batch size of 1024 check-pointing every 20 epochs.
For Normalizing Flows, we used a learning rate of $10^{-4}$, weight decay of $10^{-4}$, two layers of flows with Masked Affine Flows (Dinh et al., 2017) along with an ActNorm layer (Kingma and Dhariwal, 2018). We used a Multivariate Normal Distribution from PyTorch (Paszke et al., 2019) to build our flow model. The model was trained on 100 epochs with Adam optimizer (Kingma and Ba, 2017). All experiments used Python 3.10.
4.2 Results
Table 1 shows the performance of the open-source implementations of MolGAN and Normalizing Flows benchmarked against QM7, BBBP, Lipophilicity, PDBbind, and QM9.
Table 1: Evaluation performance of the open-source models on QM7, BBBP, Lipophilicity, PPB, QM9 datasets. The models are evaluated on the Validity of molecules generated (Val), Uniqueness(Uni), Novelty(Nov), Synthetic Accessibility Score (SAS) out of 10, and Druglikeliness(Dru). SAS is reported out of 10, whereas everything else is a %
QM7 BBBP
Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$) Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
MolGAN 28.1 32.48 100.0 8.09 25.64 54.04 5.90 100.0 6.40 38.48
Normalizing Flow 100.0 91.87 100.0 5.29 52.19 92.83 98.98 100.0 5.76 49.50
Lipophilicity PDBbind
Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$) Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
MolGAN 34.46 15.5 100.0 7.75 28.67 14.04 25.07 100.0 9.17 18.31
Normalizing Flow 100.0 99.36 100.0 6.18 49.39 100.0 98.60 100.00 5.95 50.30
QM9
Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
MolGAN 9.46 69.32 100.0 9.42 15.83
Normalizing Flow 100.0 99.15 100.0 6.58 44.55
Our results demonstrate that our implementation’s performance is comparable to existing work ( (Cao and Kipf, 2022; Kuznetsov and Polykovskiy, 2021)). We also note that all experiments reported in the paper, given the experimental setup(4.1), finish in approximately ten minutes.
5 Conclusion & Discussion
In this work, we improve DeepChem’s generative modeling tools and provide a more standardized and scalable implementation that makes generative molecular methods more accessible to scientists. Standard Machine Learning practices are built within DeepChem (e.g., checkpointing, validation, logging, etc.), which would otherwise need some form of human expertise. Benchmarks show comparable performance with existing implementations, and the tight integration with DeepChem facilitates fast future improvements. This will also allow for Reinforcement Learning (Cao and Kipf, 2022) based approaches to be incorporated easily. We will provide multi-GPU training support in future work using Distributed Data Parallelism (Li et al., 2020), Sharding & PyTorch (Paszke et al., 2019).
Impact Statement
This paper makes Generative Molecular modeling accessible to a broad spectrum of people, and while this is intended, it is not tough to modify these models to make them generate toxic and harmful molecules. However, the synthesis of novel molecules is an area that requires a lot of human expertise, and in general, is not easy to do. In a future where synthesis methods are also equally accessible, these models can be potentially dangerous, but at the present time, the positive outcomes outweigh the negative ones.
Appendix A MolGAN Pipeline
Training -
1from deepchem.models.torch_models import BasicMolGANModel as MolGAN
2from deepchem.models.optimizers import ExponentialDecay
3import deepchem as dc
4import torch.nn.functional as F
5import torch
6...
7gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
8dataset = dc.data.NumpyDataset([x.adjacency_matrix for x in features],[x.node_features for x in features])
9def iterbatches(epochs):
10 for i in range(epochs):
11 for batch in dataset.iterbatches(batch_size=gan.batch_size, pad_batches=True):
12 adjacency_tensor = F.one_hot(
13 torch.Tensor(batch[0]).to(torch.int64),
14 gan.edges).to(torch.float32)
15 node_tensor = F.one_hot(
16 torch.Tensor(batch[1]).to(torch.int64),
17 gan.nodes).to(torch.float32)
18 yield {gan.data_inputs[0]: adjacency_tensor, gan.data_inputs[1]:node_tensor}
19# train model
20gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
Inference -
1generated_data = gan.predict_gan_generator(10)
2# convert graphs to RDKitmolecules
3new_mols = feat.defeaturize(generated_data)
Appendix B NormalizngFlows Pipeline
Training -
1import deepchem as dc
2from deepchem.data import NumpyDataset
3from torch.distributions.multivariate_normal import MultivariateNormal
4from rdkit import Chem
5from dc.torch_models.nflows import *
6
7# Pass data through pipeline (mentioned above)
8...
9
10# Construct flow model
11flows = [ActNorm(latent_size)]
12nfm = NormFlow(flows, distribution, dim)
13nfm.fit(max_iterations, optimizer)
Inference -
1mols = nfm.generate(num_molecules=1000)
2valid_mols = validate_mols(mols)

### v3

@@
 Generative models for molecules have shown considerable promise for use in computational chemistry, but remain difficult to use for non-experts.
-For this reason, we introduce open-source infrastructure for easily building generative molecular models into the widely used DeepChem (Ramsundar et al., 2019) library with the aim of creating a robust and reusable molecular generation pipeline.
-In particular, we add high quality PyTorch (Paszke et al., 2019) implementations of the Molecular Generative Adversarial Networks (MolGAN)  (Cao and Kipf, 2022) and Normalizing Flows (Papamakarios et al., 2021).
-Our implementations show strong performance comparable with past work (Kuznetsov and Polykovskiy, 2021; Cao and Kipf, 2022).
-Keywords Machine Learning  $\cdot$ open-source  $\cdot$ deepchem  $\cdot$ molecule generation  $\cdot$ molecules  $\cdot$ drug discovery
+For this reason, we introduce open-source infrastructure for easily building generative molecular models into the widely used DeepChem [25] library with the aim of creating a robust and reusable molecular generation pipeline.
+In particular, we add high quality PyTorch [20] implementations of the Molecular Generative Adversarial Networks (MolGAN) [2] and Normalizing Flows [19].
+Our implementations show strong performance comparable with past work [13, 2].
 1 Introduction
-The discovery of new molecules and materials is crucial for addressing challenges in chemistry, such as treating diseases and tackling climate change (Liu et al., 2023; Sanchez and Aspuru-Guzik, 2018).
-Traditional methods rely on human expertise and are time-consuming and costly, limiting the exploration of the vast chemical space (Polishchuk et al., 2013).
-Generative models offer a promising solution using deep learning to design molecules based on desired properties, rapidly identifying diverse and optimized molecules for specific applications.
-These models vary in their approaches and have seen rapid development, with benchmarks now in place to evaluate their performance in terms of distribution learning and chemical diversity.
-Although these models are publicly available, practitioners require extensive Python and machine learning knowledge to reap their benefits.
-Thus, we introduce open-source molecular generative model infrastructure into DeepChem Ramsundar et al.
-(2019), a widely used open-source library for molecular machine learning.
-Among these models, Molecular Generative Adversarial Networks (MolGAN) (Cao and Kipf, 2022) and Normalizing Flow models (Papamakarios et al., 2021) offer two complementary approaches to the computational generation of molecular structures.
-These models, though available openly, require the practitioners to have a deep understanding of machine learning and Python, which presents significant challenges and hinders their practical adoption
-MolGAN adapts the generative adversarial network (Goodfellow et al., 2014) framework to molecular contexts, utilizing an adversarial training mechanism to facilitate the generation of novel molecular configurations by approximating the distribution of empirical molecular data.
-Normalizing Flow models use an alternative strategy through exact likelihood models, employing diffeo-morphisms to convert simple distributional forms into complex molecular distributions.
-We contribute implementations of MolGAN and Normalizing flow models in PyTorch to DeepChem and build example pipelines for training generative models.
-This adaptation broadens the accessibility of MolGAN and Normalizing Flows to the computational chemistry community.
-Our pipelines require little previous knowledge in machine learning or Python to generate molecules.
-On the other hand, the implementations are well-documented for advanced users who want to tinker around and make modifications as per their needs.
+The discovery of new molecules and materials is key to addressing challenges in chemistry, such as treating diseases and combating climate change [15, 27].
+Traditional methods, however, are time-consuming and costly, limiting the exploration of the vast chemical space [22].
+Generative models offer a deep learning-based solution, designing molecules with desired properties more efficiently.
+Despite their potential, these models typically require significant expertise in Python and machine learning.
+To address this, we introduce open-source implementations of Molecular Generative Adversarial Networks (MolGAN) [2] and Normalizing Flow models [19] in pytorch into DeepChem [25], a widely used molecular machine learning library.
+MolGAN utilizes adversarial training [7] to generate novel molecules, while Normalizing Flow models employ exact likelihood methods for molecular generation.
+Our contributions simplify their use by providing accessible pipelines that require minimal prior knowledge, while also allowing advanced users to modify the models as needed.
 2 Methods and Background
 2.1 DeepChem and Generative Molecular Models
-DeepChem (Ramsundar et al., 2019) is a highly versatile open-source Python library designed specifically for scientific machine learning and deep learning on molecular and quantum datasets (Ramsundar et al., 2021).
-Its comprehensive framework enables the solution of complex scientific problems in diverse areas such as drug discovery, energy calculations, and biotech (Wu et al., 2018).
-DeepChem’s design philosophy is centered around a systematic approach whereby scientific calculations are broken down into workflows built from underlying primitives.
-This approach enables the DeepChem library to apply to a wide range of applications.
-Notably, DeepChem has facilitated large-scale benchmarking for molecular machine learning via the MoleculeNet benchmark suite (Wu et al., 2018), protein-ligand interaction modelling (Gomes et al., 2017), generative modeling of molecules (Frey et al., 2023), and other cutting-edge scientific endeavors.
-DeepChem had older TensorFlow implementations of MolGANs and Normalizing Flows.
-However, DeepChem has been migrating to PyTorch following the adoption of PyTorch into the Linux Foundation core for long-term stability.
-In this work, we migrate the older code to PyTorch and integrate the generative framework more tightly with DeepChem’s ecosystem.
-The implementation allows users to experiment with the vast library of existing layers that can be plugged into the models to make new and creative models.
+DeepChem [25] is a versatile open-source Python library tailored for machine learning on molecular and quantum datasets [24].
+Its framework supports applications in areas such as drug discovery and biotech [32], breaking down scientific tasks into workflows built from core primitives.
+DeepChem has facilitated significant advancements, including large-scale molecular machine learning benchmarks via MoleculeNet [32], protein-ligand modeling [5], and generative molecule modeling [4].
+While older DeepChem implementations of MolGANs and Normalizing Flows used TensorFlow, we migrate these models to PyTorch, ensuring tighter integration with DeepChem’s ecosystem and broader compatibility.
+This enables users to leverage DeepChem’s extensive layer library to experiment and build new models.
 2.2 Representation of Molecules
@@
 2.2.1 One-dimensional representations
-The most common one-dimensional representation of molecules is SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988), which transforms a molecule into a sequence of characters based on predefined atom ordering rules.
+The most common one-dimensional representation of molecules is SMILES (Simplified Molecular Input Line Entry System) [31], which transforms a molecule into a sequence of characters based on predefined atom ordering rules.
 This representation enables the use of neural network architectures developed for language processing.
-For instance, previous work (Gómez-Bombarelli et al., 2018; Olivecrona et al., 2017) used recurrent neural networks as generative models to create SMILES strings.
+For instance, previous work [6, 18] used recurrent neural networks as generative models to create SMILES strings.
 However, these methods often produce invalid SMILES that cannot be converted to molecular structures due to their disregard for SMILES grammar.
-Kusner et al.
-(2017) and  Dai et al.
-(2018) attempted to address this by incorporating syntactic constraints into the recurrent neural network, yet they still frequently generated invalid SMILES strings.
-To overcome this limitation, Krenn et al.
-(2020) introduced SELFIES (Self-Referencing Embedded Strings), an improved string representation.
+To overcome this limitation, [12] introduced SELFIES (Self-Referencing Embedded Strings), an improved string representation.
 With SELFIES, a recurrent neural network can generate molecules with 100% validity, though validity here pertains to valency rules and does not guarantee molecular stability.
 2.2.2 Molecules as Graphs
-Molecules can alternatively be represented as graphs.
-Graphical representations are powerful in that they directly capture the connectivity between atoms, while in one-dimensional representations, this information must inferred by the model.
+Molecules can also be represented as graphs, where atoms are nodes and bonds are edges.
+For MolGAN, molecules are undirected graphs $G$ with edges $E$ and nodes $V$, we employ the adjacency matrix formulation, A.
+Each atom is represented by a one-hot vector, and each bond type is represented as an adjacency tensor.
+This graphical approach captures molecular connectivity directly, unlike 1D representations.
+2.3 MolGAN
+Figure 1: Model Architecture of MolGAN
+Molecular Generative Adversarial Network (MolGAN) represents a novel approach in the Generative Space for molecules by employing the GAN framework, which allows for an implicit, likelihood-free generative model which overcomes various problems (like graph matching [29] and node ordering heuristics) with previous models.
+MolGAN performs similarly to current SMILES-based approaches as well, albeit it is more susceptible to model collapse.
+We have incorporated the following loss used by the paper
+$L(\mathbf{x}^{(i)},G_{\theta}(z^{(i)});\phi)=\underbrace{-D_{\phi}(\mathbf{x}^{(i)})+D_{\phi}(G_{\theta}(z^{(i)}))}_{\text{Original WGAN Loss}}+\underbrace{\alpha(\lvert\lvert\nabla_{\hat{x}^{(i)}}D_{\phi}(\hat{x}^{(i)})\rvert\rvert-1)^{2}}_{\text{Gradient Penalty}}$ (1)
+This is an improved form of the WGAN [1] loss, where $G_{\theta}$ is the generator, $D_{\phi}$ is the discriminator, $\mathbf{x}^{(i)}\sim p_{data}(\mathbf{x})$ and $\mathbf{z}^{(i)}\sim p_{z}(\mathbf{z})$ and $\mathbf{\hat{x}}^{(i)}$ is a sampled linear combination as the following equation $\mathbf{\hat{x}}^{(i)}=\epsilon\mathbf{x}^{(i)}+(1-\epsilon)G_{\theta}(\mathbf{z}^{(i)})$ with $\epsilon\sim U(0,1)$.
+The model architecture is displayed in Figure 1.
+2.4 Normalizing Flows
+Normalizing Flows constitute a generative model that uses invertible transformations to model a probability distribution.
+This methodology enables the computation of likelihoods and the generation of samples by transforming simple base distributions (e.g., gaussian) into more complex ones through a sequence of invertible (& differentiable) transformations.
+Normalizing flows allow for direct sampling from the target distribution and thus are amazing for property-guided generation.
+The model architecture is displayed in Figure 2.
+Figure 2: Model Architecture of Normalizing Flow models
+3 Implementation
+The generative models were implemented in three main components: Layers, Base Model, and Molecule Generation Pipeline.
+We standardize (“deepchemize” [24]) molecule generation by providing BasicMolGANModel and NormalizingFlowModel which are highly flexible, allowing users to experiment with generators, discriminators, flow layers, and more.
+3.1 Layers
+MolGAN: The Discriminator in MolGAN begins with an encoder layer composed of multiple convolutional layers, followed by an aggregation layer.
+The architecture is flexible, allowing easy customization based on user requirements.
+Normalizing Flows: Normalizing Flow layers are responsible for both forward and backward computation.
+In DeepChem, the Normalizing Flows pipeline supports linear layers, which are computationally efficient compared to more complex layers like planar or autoregressive.
+These layers use linear transformations to model relationships between molecular dimensions: $\mathbf{g(x)}=\mathbf{Wx}+\mathbf{b}$ Here, $W\in\mathbb{R}^{D\times D}$ and $b\in\mathbb{R}^{D}$ are parameters, and if $W$ is invertible, the function is invertible as well.
+3.2 Base Model
+The Base Model consists of three primary components: the Generator, Discriminator, and Normalizing Flow model.
+Each component is modular and can be adjusted independently for flexibility in experimentation.
+The Generator is modeled as an MLP [17] with varying units to implicitly model a probability distribution over molecular graphs.
+It outputs continuous objects representing nodes and edges, which are transformed using the Gumbel Softmax trick [9].
+The model can also be adapted to use other methods such as the straight-through estimator [21].
+The Discriminator scores the generated molecular graph through a series of relational graph convolutional layers [28].
+These layers update node representations based on their neighbors and edge types.
+After several layers of convolution, MLPs extract the final score of the graph.
+More detailed mathematical formulations of the Generator, Discriminator, and Normalizing Flow model are mentioned in Appendix A
+3.3 Molecule Generation Pipelines
+3.3.1 MolGAN
+Figure 3: Molecule generation pipeline for MolGAN
+Training / Generating from MolGAN on custom datasets follows a very simple pipeline, which can be achieved in a few lines of code; we describe the pipeline in Figure 3, which involves extracting SMILES [31] representations and wrapping them in a DeepChem dataset that can be used to train the model.
+The full pipeline is demonstrated in B.1 with a few lines of code.
+3.3.2 Normalizing Flows
+Figure 4: Molecule generation pipeline for Normalizing flows
+Analogous to the MolGAN pipeline, the Normalizing Flows pipeline (Figure 4) expects a SELFIES [12] string as input.
+There are added steps for Dequantization (i.e., adding noise in [0, 1) to every input) The full pipeline is demonstrated in B.2 with a few lines of code.
+4 Experiments
+4.1 Datasets
+We use the following publicly available datasets: QM7 [26], BBBP [16], Lipophilicity [8], PPB [30], and QM9 [23].
+QM7 is a subset of GDB-13, consisting of molecules with up to 7 heavy atoms (C, N, O, and S).
+The BBBP dataset includes around 2,000 molecules with binary labels related to blood-brain barrier permeability.
+The Lipophilicity dataset, sourced from ChEMBL, contains 4,200 compounds.
+The PPB dataset, curated from PubChem BioAssay by the Maximum Unbiased Validation (MUV) group, includes around 11,000 compounds.
+QM9 is a subset of GDB-17, comprising 134,000 organic molecules with up to 9 heavy atoms.
+All datasets were obtained through MoleculeNet [32].
+For nomalizing flows we utilize the full dataset for all experiments, however, following [2], for MolGAN, we only use molecules that only have C, N, O and F in them.
+The final number of valid molecules in every dataset are shown in 1.
+Dataset Number of Samples
+BBBP 1631
+PPB 1291
+QM7 5470
+QM9 105984
+Lipophilicity 3360
+Table 1: Number of samples in each dataset.
+4.2 Experimental Setup
+For MolGAN, we use four edge types (single, double, triple, and no-bond), five node types (C, N, O, F, and PAD) similar to [2], whilst our discriminator discriminates among molecules of all lengths, our generator is restricted to small molecule generation, limited to 9 vertices.
+To allow this, we would set the maximum number of allowed atoms to the maximum in a dataset.
+We report mean results across 10 seeds, each involving generation of 6400 molecules, as in [2].
+For Normalizing Flows, we used two layers of flows with Masked Affine Flows [3] along with an ActNorm layer [11].
+We used a Multivariate Normal Distribution from PyTorch [20] to build our flow model.
+All hyper-parameters are listed in C.
+All experiments used Python 3.10.
+4.3 Results
+Dataset Model Val($\uparrow$) Uni($\uparrow$) Nov($\uparrow$) SAS($\downarrow$) Dru($\uparrow$)
+QM7 MolGAN $92.1$ $4.18$ $100.0$ $2.11$ $70.45$
+Norm Flow $100.0$ $91.87$ $100.0$ $5.29$ $52.19$
+QM9 MolGAN $89.67$ $4.46$ $100.0$ $2.83$ $66.03$
+Norm Flow $100.0$ $99.15$ $100.0$ $6.58$ $44.55$
+Original Implementation [2] $87.7$ $2.9$ $97.7$ —-* —-*
+BBBP MolGAN $57.18$ $0.28$ $100.0$ $5.24$ $55.42$
+Norm Flow $92.83$ $98.98$ $100.0$ $5.76$ $49.50$
+Lipophilicity MolGAN $49.57$ $0.08$ $100.0$ $5.75$ $53.58$
+Norm Flow $100.0$ $99.36$ $100.0$ $6.18$ $49.39$
+PPB MolGAN $14.04$ $25.07$ $100.0$ $9.17$ $18.31$
+Norm Flow $100.0$ $98.60$ $100.0$ $5.95$ $50.30$
+Table 2: Evaluation performance of the open-source models on QM7, BBBP, Lipophilicity, PPB, QM9 datasets.
+The models are evaluated on the Validity of molecules generated (Val), Uniqueness (Uni), Novelty (Nov), Synthetic Accessibility Score (SAS) out of 10, and Druglikeliness (Dru).
+SAS is reported out of 10, whereas everything else is a %.
+(* - not reported in the paper)
+Our results 2 demonstrate that our implementation’s performance is comparable to existing work ( [2, 13]).
+For a fair comparison, we only compare our implementation to the non-RL results from [2], also [2] only train and evaluate on QM9.
+5 Conclusion & Discussion
+In this work, we improve DeepChem’s generative modeling tools and provide a more standardized and scalable implementation that makes generative molecular methods more accessible to scientists.
+Standard Machine Learning practices are built within DeepChem (e.g., checkpointing, validation, logging, etc.), which would otherwise need some form of human expertise.
+Benchmarks show comparable performance with existing implementations, and the tight integration with DeepChem facilitates fast future improvements.
+This will also allow for Reinforcement Learning [2] based approaches to be incorporated easily.
+We observed high variance for datasets where the average number of atoms in a molecule was much larger than 9, and leave this for future work.
+We will provide multi-GPU training support in future work using Distributed Data Parallelism [14], Sharding & PyTorch [20].
+Impact Statement
+This paper makes Generative Molecular modeling accessible to a broad spectrum of people, and while this is intended, it is not tough to modify these models to make them generate toxic and harmful molecules.
+However, the synthesis of novel molecules is an area that requires a lot of human expertise, and in general, is not easy to do.
+In a future where synthesis methods are also equally accessible, these models can be potentially dangerous, but at the present time, the positive outcomes outweigh the negative ones.
+Appendix
+Appendix A Mathematical Formulations
+A.1 Molecules as Graphs
 For MolGAN, each molecule can be represented as an undirected Graph $G$ with a set of edges $E$ and nodes $V$.
@@
 $\mathbf{A}_{i,j}\in\mathbb{R}^{Y}$ is a one-hot vector indicating the type of edge between $i$ and $j$.
-2.3 MolGAN
-Molecular Generative Adversarial Network (MolGAN) represents a novel approach in the Generative Space for molecules by employing the GAN framework, which allows for an implicit, likelihood-free generative model which overcomes various problems (like graph matching (Simonovsky and Komodakis, 2018) and node ordering heuristics) with previous models.
-MolGAN performs similarly to current SMILES-based approaches as well, albeit it is more susceptible to model collapse.
-We have incorporated the following loss used by the paper
-$L(\mathbf{x}^{(i)},G_{\theta}(z^{(i)});\phi)=\underbrace{-D_{\phi}(\mathbf{x}^{(i)})+D_{\phi}(G_{\theta}(z^{(i)}))}_{\text{Original WGAN Loss}}+\\ \underbrace{\alpha(\lvert\lvert\nabla_{\hat{x}^{(i)}}D_{\phi}(\hat{x}^{(i)})\rvert\rvert-1)^{2}}_{\text{Gradient Penalty}}$ (1)
-This is an improved from of the WGAN (Arjovsky et al., 2017) loss, where $G_{\theta}$ is the generator, $D_{\phi}$ is the discriminator, $\mathbf{x}^{(i)}\sim p_{data}(\mathbf{x})$ and $\mathbf{z}^{(i)}\sim p_{z}(\mathbf{z})$ and $\mathbf{\hat{x}}^{(i)}$ is a sampled linear combination as the following equation $\mathbf{\hat{x}}^{(i)}=\epsilon\mathbf{x}^{(i)}+(1-\epsilon)G_{\theta}(\mathbf{z}^{(i)})$ with $\epsilon\sim U(0,1)$.
-The model architecture is displayed in Figure 1.
-Figure 1: Model Architecture of MolGAN
-2.4 Normalizing Flows
-Normalizing Flows constitute a generative model that uses invertible transformations to model a probability distribution.
-This methodology enables the computation of likelihoods and the generation of samples by transforming simple base distributions (e.g., gaussian) into more complex ones through a sequence of invertible (& differentiable) transformations.
-The precision and control afforded by Normalizing Flows make them particularly advantageous for generating molecules with highly specific attributes.
-Normalizing flows also take advantage of equivariance (through invariance) in the Graph’s probability distribution, thus providing a complementary strength to adversarial approaches such as MolGAN in the generative modeling landscape.
-The model architecture is displayed in Figure 2.
-Figure 2: Model Architecture of Normalizing Flow models
-2.5 Standardizing Molecule Generation Workflows
-We aim to standardize (“deepchemize”  (Ramsundar et al., 2021)) the molecule generation by MolGAN and Normalizing Flow model by implementing the models using the workflow structure implemented by other DeepChem models.
-With this standardization, we aim to experimenting with new generative models simpler and enable better molecule generation capabilities.
-BasicMolGANModel (the MolGAN implementation in DeepChem), can be used with different discriminators and generators, thus providing a flexible framework for molecule generation.
-This flexibility of the BasicMolGANModel can be used to incorporate more complex architectures of generators and discriminators to achieve better global density information to approximate a more accurate exchange-correlation functional.
-We allow for a “plug-and-play" style of training, inference, and generation with respect to the MPNN used, optimizer, number of MLP layers in the generator, and bond/node types.
-This can be used in a creatively to combine the strengths of multiple architectures.
-NormalizingFlowModel (the Normalizing Flow implementation in DeepChem) can be used with different flow layers with customized bijective transformations.
-This allows end users to incorporate more complex transformations and modeling capabilities into the framework.
-The interoperable nature of DeepChem’s layers enables further innovation.
-Overall, these abstractions simplify and ease usage and experimentation to achieve state-of-the-art molecule generation.
-We have also created pedagogical tutorials for all layers and networks introduced to provide a streamlined pipeline for end-users, as well as, enable fast experimentation as the mutable parameters are clearly demarcated.
-3 Implementation
-The Generative models were split and implemented as the following major components: Layers, Base Model, and Molecule Generation Pipeline. We briefly also describe the DeepChem Datasets used.
-3.1 Datasets
-Our work utilizes three publicly available datasets: QM7 (Rupp et al., 2012), BBBP (Martins et al., 2012), Lipophilicity (Hersey, 2015), PPB (Wang et al., 2005) and QM9 (Ramakrishnan et al., 2014).
-QM7 is a subset of GDB-13 (a database of nearly 1 billion stable and synthetically accessible organic molecules) containing up to 7 heavy atoms C, N, O, and S.
-The blood-brain barrier penetration (BBBP) dataset encompasses approximately 2,000 molecules with binary labels on their permeability properties.
-The lipophilicity dataset, curated from ChEMBL database, contains 4200 compounds.
-The PPB dataset is from the Maximum Unbiased Validation (MUV) group, it is a benchmark dataset selected from PubChem BioAssay by applying a refined nearest neighbor analysis.
-It contains around eleven thousand compounds.
-QM9 is a subset of the GDB-17 database (a database of nearly 166 billion stable and synthetically accessible organic molecules), comprising 134 thousand stable organic molecules with up to 9 heavy atoms.
-All the mentioned datasets are downloaded through MoleculeNet (Wu et al., 2018).
-3.2 Layers
-MolGAN: The first layer in the Discriminator of a MolGAN model is an encoder layer, which consists of multiple convolutional layers followed by an aggregation layer.
-We provide an architecture that can be varied and overridden easily based on the end user’s requirements.
-Normalizing Flows: Layers in Normalizing Flows are used to perform forward and backward computation.
-DeepChem Normalizing Flows pipeline supports linear layers, which offer low computational costs compared with other layers (e.g.
-planar, autoregressive).
-Linear transformations exhibit relationships through molecular data dimensions.
-$\mathbf{g(x)}=\mathbf{Wx}+\mathbf{b}$
-Where $W\in\mathbb{R}^{D\times D}$ and $b\in\mathbb{R}^{D}$ are parameters. So, if W is an invertible matrix, the function is invertible.
-3.3 Base Model
-The Base Model has three distinct components, which can be modified individually to enable flexibility.
-[…]
+A.2 Generator
+For any latent variable
 $z\sim\mathcal{N}(\mathbf{0,I})$
-the generator $\mathcal{G}_{\theta}(z)$ outputs two continuous and dense objects,
+the generator $\mathcal{G}_{\theta}(z)$ outputs two continuous objects:
 $\mathbf{X}\in\mathbb{R}^{N\times D},\mathbf{A}\in\mathbb{R}^{N\times N\times D}$
-Over which we perform categorical sampling using the Gumbel Softmax trick (Jang et al., 2017) to obtain $\mathbf{\tilde{X}}$ and $\mathbf{\tilde{A}}$ as -
+Using the Gumbel Softmax trick [9], categorical sampling is performed as:
 $\mathbf{\tilde{X}}$ $=\mathbf{X}+\text{Gumbel}(\mu=0,\beta=1)$ (2)
 $\mathbf{\tilde{A}}$ $=\mathbf{A}_{ijy}+\text{Gumbel}(\mu=0,\beta=1)$ (3)
-This method can be overridden to use a straight-through estimator (Paulus et al., 2020) or can also allow for $\mathbf{\tilde{X}}=\mathbf{X}$ and $\mathbf{\tilde{A}}=\mathbf{A}$.
-Discriminator - The discriminator is tasked with scoring the graph produced. This involves a series of graph-convolutional or message-passing layers followed by MLPs; the equations are as follows:
+Alternatively, $\mathbf{\tilde{X}}$ and $\mathbf{\tilde{A}}$ can remain as $\mathbf{X}$ and $\mathbf{A}$.
+A.3 Discriminator
+The discriminator updates node representations through relational graph convolution as:
 $\boldsymbol{h}_{i}^{\prime(\ell+1)}$ $=f_{s}^{(\ell)}\left(\boldsymbol{h}_{i}^{(\ell)},\boldsymbol{x}_{i}\right)+\sum_{j=1}^{N}\sum_{y=1}^{Y}\frac{\tilde{\boldsymbol{A}}_{ijy}}{\left|\mathcal{N}_{i}\right|}f_{y}^{(\ell)}\left(\boldsymbol{h}_{j}^{(\ell)},\boldsymbol{x}_{j}\right)$ (4)
 $\boldsymbol{h}_{i}^{(\ell+1)}$ $=\tanh\left(\boldsymbol{h}_{i}^{\prime(\ell+1)}\right)$ (5)
-We employ Relational GCN (Schlichtkrull et al., 2017) layers, which allow us to model the graph structure with multiple edge types, this shows the update of node representations per layer.
-Here, $h_{i}^{(l)}$ represents the signal of node $i$ at layer $l$ and $f_{s}^{(l)}$ represents a self-connection linear transformation.
-$\mathcal{N}_{i}$ represents the neighbours of node $i$.
+where $\boldsymbol{h}_{i}^{(l)}$ represents the signal at node $i$ and layer $l$, and $f_{s}^{(l)}$ is a self-connection transformation.
+After several layers of graph convolution, the graph score is calculated as:
 $\boldsymbol{h}_{\mathcal{G}}^{\prime}$ $=\sum_{v\in\mathcal{V}}\sigma\left(i\left(\boldsymbol{h}_{v}^{(L)},\boldsymbol{x}_{v}\right)\right)\odot\tanh\left(j\left(\boldsymbol{h}_{v}^{(L)},\boldsymbol{x}_{v}\right)\right)$ (6)
 $\boldsymbol{h}_{\mathcal{G}}$ $=\tanh(\boldsymbol{h}_{\mathcal{G}}^{\prime})$ (7)
-After a few layers of graph convolution, we use two MLPs $i$ and $j$ similar to  (Li et al., 2017) to extract the overall graph score.
-$\sigma$ stands for the sigmoid function $\sigma=\frac{1}{1+e^{-x}}$.
-Normalizing Flow model starts with a base distribution $p_{z}$ and performs a series of layer transformations to estimate a density function $p_{x}$, i.e.,
-$x=L(z)$
-where $z\sim p_{z}$, importantly, these transformations must be invertible; and so,
+where $\sigma$ is the sigmoid function $\sigma=\frac{1}{1+e^{-x}}$.
+A.4 Normalizing Flow Model
+The Normalizing Flow model starts with a base distribution $p_{z}$ and transforms it into a density function $p_{x}$.
+The transformation $x=L(z)$ is invertible, and the probability density is updated as:
 $\rho_{X}(x)=\rho_{Z}(z)\cdot|det(J_{L}(z))|^{-1}$
-Jacobian determinant, the second term on the right side of the equation, updates the direction of the flow based on linear transformations.
-3.4 Molecule Generation Pipelines
-3.4.1 MolGAN
-Figure 3: Molecule generation pipeline for MolGAN
-Training / Generating from MolGAN on custom datasets follows a very simple pipeline, which can be achieved in a few lines of code; we describe the pipeline in Figure 3, which involves extracting SMILES (Weininger, 1988) representations and wrapping them in a DeepChem dataset that can be used to train the model.
-gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
-gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
-The full pipeline is in Appendix A.
-3.4.2 Normalizing Flows
-Figure 4: Molecule generation pipeline for Normalizing flows
-Analogous to the MolGAN pipeline, the Normalizing Flows pipeline (Figure 4) expects a SELFIES (Krenn et al., 2020) string as input.
-There are added steps for Dequantization (i.e., adding noise in [0, 1) to every input) and test-train-valid splitter (unlike MolGAN, where you train on the full dataset).
-1
-2nfm = NormFlow(flows, distribution, dim)
-3
-4nfm.fit(max_iterations, optimizer)
-5
-The full pipeline is in Appendix B.
-4 Experimental Results
-4.1 Experimental Setup
-We report results averaged across five runs of 1000 molecule generations, matching Cao and Kipf (2022).
-For MolGAN evaluation, we allowed four edge types (single, double, triple, no-bond, and aromatic), five node types (C, N, O, F, and PAD), and a maximum of nine vertices per graph; this is borrowed from the original implementation along with a learning rate of $10^{-3}$.
-However, we use an embedding dimension of 100, no dropout (Srivastava et al., 2014), and a batch size of 1024 check-pointing every 20 epochs.
-For Normalizing Flows, we used a learning rate of $10^{-4}$, weight decay of $10^{-4}$, two layers of flows with Masked Affine Flows (Dinh et al., 2017) along with an ActNorm layer (Kingma and Dhariwal, 2018).
-We used a Multivariate Normal Distribution from PyTorch (Paszke et al., 2019) to build our flow model.
-The model was trained on 100 epochs with Adam optimizer (Kingma and Ba, 2017).
-All experiments used Python 3.10.
-4.2 Results
-Table 1 shows the performance of the open-source implementations of MolGAN and Normalizing Flows benchmarked against QM7, BBBP, Lipophilicity, PDBbind, and QM9.
-Table 1: Evaluation performance of the open-source models on QM7, BBBP, Lipophilicity, PPB, QM9 datasets.
-The models are evaluated on the Validity of molecules generated (Val), Uniqueness(Uni), Novelty(Nov), Synthetic Accessibility Score (SAS) out of 10, and Druglikeliness(Dru).
-SAS is reported out of 10, whereas everything else is a %
-QM7 BBBP
-Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$) Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
-MolGAN 28.1 32.48 100.0 8.09 25.64 54.04 5.90 100.0 6.40 38.48
-Normalizing Flow 100.0 91.87 100.0 5.29 52.19 92.83 98.98 100.0 5.76 49.50
-Lipophilicity PDBbind
-Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$) Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
-MolGAN 34.46 15.5 100.0 7.75 28.67 14.04 25.07 100.0 9.17 18.31
-Normalizing Flow 100.0 99.36 100.0 6.18 49.39 100.0 98.60 100.00 5.95 50.30
-QM9
-Val ($\uparrow$) Uni ($\uparrow$) Nov ($\uparrow$) SAS ($\downarrow$) Dru ($\uparrow$)
-MolGAN 9.46 69.32 100.0 9.42 15.83
-Normalizing Flow 100.0 99.15 100.0 6.58 44.55
-Our results demonstrate that our implementation’s performance is comparable to existing work ( (Cao and Kipf, 2022; Kuznetsov and Polykovskiy, 2021)).
-We also note that all experiments reported in the paper, given the experimental setup(4.1), finish in approximately ten minutes.
-5 Conclusion & Discussion
-In this work, we improve DeepChem’s generative modeling tools and provide a more standardized and scalable implementation that makes generative molecular methods more accessible to scientists.
-Standard Machine Learning practices are built within DeepChem (e.g., checkpointing, validation, logging, etc.), which would otherwise need some form of human expertise.
-Benchmarks show comparable performance with existing implementations, and the tight integration with DeepChem facilitates fast future improvements.
-This will also allow for Reinforcement Learning (Cao and Kipf, 2022) based approaches to be incorporated easily.
-We will provide multi-GPU training support in future work using Distributed Data Parallelism (Li et al., 2020), Sharding & PyTorch (Paszke et al., 2019).
-Impact Statement
-This paper makes Generative Molecular modeling accessible to a broad spectrum of people, and while this is intended, it is not tough to modify these models to make them generate toxic and harmful molecules.
-However, the synthesis of novel molecules is an area that requires a lot of human expertise, and in general, is not easy to do.
-In a future where synthesis methods are also equally accessible, these models can be potentially dangerous, but at the present time, the positive outcomes outweigh the negative ones.
-Appendix A MolGAN Pipeline
+where $J_{L}(z)$ is the Jacobian determinant of the transformation.
+Appendix B Code Snippets
+B.1 MolGAN Pipeline
 Training -
-1from deepchem.models.torch_models import BasicMolGANModel as MolGAN
-2from deepchem.models.optimizers import ExponentialDecay
-3import deepchem as dc
-4import torch.nn.functional as F
-5import torch
-6...
-7gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
-8dataset = dc.data.NumpyDataset([x.adjacency_matrix for x in features],[x.node_features for x in features])
-9def iterbatches(epochs):
+1 from deepchem.models.torch_models import BasicMolGANModel as MolGAN
+2 from deepchem.models.optimizers import ExponentialDecay
+3 import deepchem as dc
+4 import torch.nn.functional as F
+5 import torch
+6 ...
+7 gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
+8 dataset = dc.data.NumpyDataset([x.adjacency_matrix for x in features],[x.node_features for x in features])
+9 def iterbatches(epochs):
 10 for i in range(epochs):
@@
 18 yield {gan.data_inputs[0]: adjacency_tensor, gan.data_inputs[1]:node_tensor}
-19# train model
-20gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
+19 # train model
+20 gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
 Inference -
-1generated_data = gan.predict_gan_generator(10)
-2# convert graphs to RDKitmolecules
-3new_mols = feat.defeaturize(generated_data)
-Appendix B NormalizngFlows Pipeline
+1 generated_data = gan.predict_gan_generator(10)
+2 # convert graphs to RDKitmolecules
+3 new_mols = feat.defeaturize(generated_data)
+B.2 NormalizngFlows Pipeline
 Training -
-1import deepchem as dc
-2from deepchem.data import NumpyDataset
-3from torch.distributions.multivariate_normal import MultivariateNormal
-4from rdkit import Chem
-5from dc.torch_models.nflows import *
+1 import deepchem as dc
+2 from deepchem.data import NumpyDataset
+3 from torch.distributions.multivariate_normal import MultivariateNormal
+4 from rdkit import Chem
+5 from dc.torch_models.nflows import *
 6
-7# Pass data through pipeline (mentioned above)
-8...
+7 # Pass data through pipeline (mentioned above)
+8 ...
 9
-10# Construct flow model
-11flows = [ActNorm(latent_size)]
-12nfm = NormFlow(flows, distribution, dim)
-13nfm.fit(max_iterations, optimizer)
+10 # Construct flow model
+11 flows = [ActNorm(latent_size)]
+12 nfm = NormFlow(flows, distribution, dim)
+13 nfm.fit(max_iterations, optimizer)
 Inference -
-1mols = nfm.generate(num_molecules=1000)
-2valid_mols = validate_mols(mols)
+1 mols = nfm.generate(num_molecules=1000)
+2 valid_mols = validate_mols(mols)
+Appendix C Hyperparameters
+For our MolGAN implementation, similar to [2], we vary the given hyperparameters, and pick the best for each dataset prioritising Validity of generations.
+Except for QM9, we train each model for 30 epochs on the whole dataset.
+QM9 is trained similar to  [2], i.e., 300 epochs on a random 5k subset, we also perform early stopping if the uniqueness goes below 2%.
+Apart from this we fix Generator Steps to 0.2, i.e., one step every 5 discriminator steps, a batch size of 32, learning rate of $41e-$ 4, and an embedding dimension of 32.
+Hyperparameter Values
+Dropout Rates $\{0.0,0.1,0.25\}$
+Sampling Mode Straight-through, Gumbel, SoftMax
+Table 3: Hyperparameter settings for MolGAN.
+For the Normalizing Flows implementation, we use a learning rate of 1e-4, and an equal weight decay.
+We train on a batch size of 1024 for 100 epochs.
+Both the models were trained on with Adam optimizer [10].